POPULATION TRANFER SPECTROSCOPY: NEW PROBES OF THE DYNAMICS OF FLEXIBLE BIOMOLECULES

\underline{BRIAN C. DIAN}, JASPER R. CLARKSON , GINA M. FLORIO, ASIER LONGARTE, and TIM-OTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907-1393.

Infrared Population Transfer Spectroscopy (IRPTS) and Hole-Filling Spectroscopy (HFS) have recently been developed by our group as a means to study gas-phase conformational dynamics of flexible biomolecules. These methods utilize selective infrared excitation of single conformations of the molecules of interest in the early portions of a gas-phase expansion, followed by collisional re-cooling of the excited population into its conformational minima for subsequent conformationspecific detection. These techniques can serve as a powerful probe to study both collisional and conformational dynamics of flexible biomolecules and their associated water clusters in a supersonic expansion. We have recently demonstrated the ability to determine Population Transfer Quantum Yields for dipeptide analogs of tryptophan and the hormone melatonin; molecules for which a large number of conformational degrees of freedom exist. In addition, we are able to determine semiquantitative information about barriers to intramolecular conformational isomerization, quantitative information about the fractional populations of the conformers in the absence of infrared excitation and the conformational consequences of X-(H_{2}O)_{n} dissociation dynamics. Application of the aforementioned techniques to a selection of mimetic peptide oligomers will be discussed.